A SHORT STORY

by A. U. Thor

Once upon a time, in a distant galaxy called Ööç, there lived a computer named R. J. Drofnats. Mr. Drofnats—or “R. J.,” as he preferred to be called— was happiest when he was at work typesetting

beautiful documents.